# Surrogate-Assisted Targeted Learning for Delayed Outcomes under Administrative Censoring

### 11 Mar 2026 v1

@@ Abstract
Delayed primary outcomes and administratively censored follow-up create a general semiparametric estimation problem: the target causal functional depends on an endpoint observed only for a shrinking subset of units at analysis time, while earlier surrogate measurements remain widely available. In such settings, inverse-probability-weighted estimators can become unstable as observation probabilities approach the positivity boundary, and complete-case model-based analyses can be highly sensitive to outcome-model specification.
We develop a surrogate-assisted targeted minimum loss estimator for this nested causal functional. Identification proceeds through a surrogate-bridge representation that integrates an observed-outcome regression over the conditional surrogate distribution, thereby avoiding inverse observation weights in the target parameter itself. We show that the estimator is asymptotically linear and doubly robust (in the sense that first-order bias vanishes when either nuisance component is consistently estimated), and we characterize two structural features of the problem: under surrogate-mediated missing at random, the censoring mechanism contributes no separate tangent-space component to the efficient influence function; and for nested bridge functionals, a one-step debiased machine-learning construction leaves a second-order cross-product remainder involving the conditional surrogate law. The proposed two-stage targeting step removes this term without requiring direct estimation of that law.
Simulation studies demonstrate stable finite-sample performance under substantial administrative censoring, and a design-calibrated analysis based on the Washington State EPT study illustrates the method in a realistic stepped-wedge cluster-randomized setting.
@@ 1 Introduction
Many modern studies measure the primary endpoint only after substantial delay, while shorter-term surrogate outcomes are available much earlier in follow-up. When analysis occurs before all primary outcomes mature, the resulting observed-data structure poses a distinctive semiparametric problem: the estimand depends on a delayed endpoint observed only for a subset of units whose observation probability may approach the positivity boundary, yet a surrogate remains broadly observed across the design space. In such settings, estimators that rely directly on inverse observation weights can become unstable as observation probabilities become small, whereas complete-case model-based analyses may inherit strong sensitivity to outcome-model specification.
This paper studies that problem through a surrogate-bridge formulation for causal effect estimation under administrative censoring. The key idea is to identify the target functional — the average treatment effect $\Psi(P_{0})=\mathbb{E}[Y(1)-Y(0)]$ — by integrating an observed-outcome regression over the conditional surrogate law, rather than by placing the inverse observation probability directly in the target parameter. This perspective leads naturally to a targeted learning estimator and exposes new structure in the corresponding semiparametric theory.
A stepped-wedge cluster randomized trial (SW-CRT) provides a particularly sharp motivating example. Because treatment rollout is tied to calendar time, late-crossing clusters may have little opportunity to realize the delayed endpoint before study closure. But the underlying statistical issue is broader than stepped-wedge designs: whenever an early surrogate is observed widely and a primary endpoint is administratively censored at analysis time, the analyst faces the same nested surrogate-bridge problem. We therefore use the stepped-wedge setting as a concrete and practically important instance of a more general semiparametric estimation problem.
@@
Parametric mixed models (Hussey and Hughes, 2007; Hemming and Taljaard, 2020; Hughes et al., 2020) restrict attention to the complete-case sub-cohort $\{\Delta_{ijt}=1\}$ and require correct specification of both the secular time trend and the missing-data mechanism. IPCW estimators (Robins et al., 1994; Bang and Robins, 2005) place $g_{\Delta}^{-1}$ directly in the estimating equation; even moderate concentration of $g_{\Delta}$ near zero for late-crossing clusters produces severe variance inflation, and in the near-boundary regime these estimators may fail to admit regular asymptotic approximations.
@@
This paper makes three methodological contributions.
First, we identify a class of causal functionals under surrogate-mediated administrative censoring through a nested bridge representation (Theorem 1) that avoids inverse observation weights in the target parameter. The surrogate-bridge G-computation formula identifies $\Psi(P_{0})$ via nested conditional expectations over the observed surrogate distribution, replacing the unstable $g_{\Delta}^{-1}$ with a support positivity condition on the complete-case outcome regression.
Second, we characterize the efficient-score structure of this problem (Section 3): under surrogate-mediated missing at random, the censoring mechanism contributes no separate tangent-space component to the efficient influence function (Lemma 1), so estimating $g_{\Delta}$ does not reduce the efficiency bound; and under clustered dependence, valid inference requires aggregation of influence contributions at the cluster level rather than averaging (Lemma 2).
Third, we show that for this nested functional, a standard one-step debiased machine-learning construction generally leaves a second-order cross-product remainder $R_{SY}$ involving the conditional surrogate law $f_{S}$ (Proposition 2); the proposed two-stage targeted minimum loss procedure removes that term without requiring direct estimation of $f_{S}$, achieving $\sqrt{J}$-consistency under a product-rate condition on the remaining nuisance estimators (Theorem 2).
@@
Our work sits at the intersection of semiparametric missing-data theory, targeted learning, and modern debiased machine learning. Classical AIPW theory (Robins et al., 1994; Bickel et al., 1993) characterizes efficient estimation with missing outcomes via tangent space decomposition, but does not address the nested bridge structure considered here. Existing SW-CRT methodology (Hussey and Hughes, 2007; Hemming and Taljaard, 2020; Kasza et al., 2019; Hughes et al., 2020) largely remains centered on parametric mixed models and does not accommodate the present surrogate-assisted functional. The surrogate outcome literature (Gilbert et al., 2008; Frangakis and Rubin, 2002) addresses a complementary problem — evaluating treatment effects when the primary outcome is rare, not delayed — and their identification strategies do not extend to the near-boundary censoring regime. Luedtke et al. (2017) construct the closest prior estimator: a semiparametric surrogate-index estimator for i.i.d. data whose identification formula is structurally related to ours. However, their asymptotic linearity argument requires a uniform positivity lower bound on $g_{\Delta}$, which the SW-CRT’s calendar-driven censoring structure violates for late-crossing clusters; the present surrogate-bridge approach does not place $g_{\Delta}^{-1}$ in the target parameter and therefore does not require this condition. Cluster-level targeted learning (Balzer et al., 2016, 2021) establishes the importance of clusterwise influence aggregation for valid sandwich inference in group-randomized trials, but without the bridge-based identification and nested targeting problem studied here. The debiased machine learning framework (Chernozhukov et al., 2018) provides general product-rate guarantees via sample splitting, yet in the present setting a one-step construction leaves a nested cross-product remainder $R_{SY}$ involving the conditional surrogate law $f_{S}$ that cross-fitting does not eliminate. The contribution of this paper is to isolate that remainder explicitly (Proposition 2) and show how a second targeting step absorbs it without estimating $f_{S}$. [^1] See Kennedy (2022) for a comprehensive review of debiased machine learning. Proposition 2 gives the formal derivation.
@@
Section 2 defines the observed data structure and establishes non-parametric identification via the surrogate-bridge representation. Section 3 develops the semiparametric theory: the EIC decomposition, the vanishing censoring-mechanism component (Lemma 1), the cluster-level summation rule (Lemma 2), and the nested cross-product remainder that makes two-stage targeting necessary (Proposition 2). Section 4 constructs the SA-TMLE as the direct solution to the theoretical constraints identified in Section 3. Section 5 establishes asymptotic linearity, double robustness, and the Berry–Esseen finite-sample coverage bound. Section 6 reports Monte Carlo experiments across three scenario blocks. Section 7 provides a brief design-calibrated illustration using the Washington State EPT trial.
Although our motivating application is a stepped-wedge cluster trial, the framework is not tied to that design. The same statistical structure arises whenever a delayed primary outcome is observed only for a subset of units at analysis time, an earlier surrogate is broadly available, and identification proceeds through a nested bridge representation. The paper is best understood not as a stepped-wedge method but as a general semiparametric framework for nested surrogate-bridge inference under administrative censoring, with stepped-wedge trials providing a particularly consequential use case.
Table 1: When to use each estimator: a practical comparison for stepped-wedge trials with delayed outcomes.
GLMM IPCW/AIPW SA-TMLE (proposed)
Target parameter Conditional treatment effect (model-based) Population ATE Population ATE
Key assumptions Correct outcome model specification (secular trend, random effects) Observation positivity: $g_{\Delta}^{-1}$ must be well-behaved Support positivity for $E[Y\mid S,A,W,t]$; surrogate-mediated MAR
Handles heavy censoring? No: complete-case analysis drops late-crossing clusters entirely Poorly: variance inflates as $g_{\Delta}\to 0$ Yes: bypasses $g_{\Delta}^{-1}$ via surrogate bridge
Requires surrogate? No No Yes
Double robustness? No Yes (but requires bounded $g_{\Delta}^{-1}$) Yes (asymptotically; finite-sample behavior may be asymmetric under nuisance misspecification)
Nonparametric nuisance? No (parametric outcome model) Yes (machine learning compatible) Yes (machine learning compatible)
Best suited for Low censoring, well-understood secular trends, no surrogate available Moderate censoring with good observation-positivity overlap Substantial administrative censoring with an available early surrogate
@@ 2 Data Structure and Identification
We first describe the observed-data structure and identification strategy in the stepped-wedge cluster-randomized setting, which serves as our running example of a broader class of surrogate-assisted problems with administratively censored delayed outcomes.
@@
Three structural features deserve emphasis. First, $A_{jt}$ is deterministic given $\tau_{j}$; the treatment mechanism is known by design and does not require data-adaptive estimation (see Remark 4). Second, the cluster random effect $b_{j}$ enters both $f_{S}$ and $f_{Y}$, inducing the ICC among all observations within cluster $j$ (Remark 1). Third — and most critically — $Y$ does not appear in $f_{\Delta}$ and $U_{\Delta}$ does not appear in $f_{Y}$. This is the structural encoding of Assumption 3: conditional on $S$, the censoring indicator $\Delta$ is independent of the outcome $Y$. In DAG terms, there is no directed edge $Y_{ijt}\to\Delta_{ijt}$ (Figure 1).
@@
The shared cluster random effect $b_{j}\overset{\mathrm{i.i.d.}}{\sim}\mathcal{N}(0,\sigma^{2}_{b})$ enters both $f_{S}$ and $f_{Y}$, inducing ICC $\sigma^{2}_{b}/(\sigma^{2}_{b}+\sigma^{2}_{\varepsilon})$ among individuals within a cluster. The two methodological consequences are: (i) inference must use the cluster-robust sandwich estimator of Section 3; and (ii) Super Learner cross-validation must be performed at the cluster level (Supplement, Section S3). The TMLE estimator itself makes no parametric assumption about $b_{j}$; the Gaussian model governs the simulation DGP only.
@@ 2.1 Potential Outcomes and the Causal Target
For intervention value $a\in\{0,1\}$, let $S_{ijt}(a)$ and $Y_{ijt}(a)$ denote the potential outcomes obtained by the intervention $\mathrm{do}(A_{jt}=a)$ in the SCM. The causal Average Treatment Effect (ATE) is:
$\Psi(P_{0})\;=\;\mathbb{E}\bigl[Y_{ijt}(1)-Y_{ijt}(0)\bigr],$ (7)
where the expectation is over the marginal distribution of $(W_{ijt},b_{j},t)$ in the trial. The identification of $\Psi(P_{0})$ from $P_{0}$ is the subject of Section 2.2.
@@ 2.2 Non-Parametric Identification
The central identification insight is that the target causal effect can be represented through a surrogate bridge: one first learns the observed-outcome regression among units with realized primary outcomes and then integrates that regression over the treatment-specific surrogate distribution. This representation identifies the causal estimand without placing inverse observation weights in the target functional itself, which is precisely what makes it attractive in regimes where observation probabilities become small.
We establish the following assumptions to formalize this representation.
@@
Conditional on the cluster history, treatment is independent of potential outcomes:
$\bigl(Y(0),Y(1),S(0),S(1)\bigr)\;\perp\;A_{jt}\;\big|\;W_{ijt},\,t.$ (8)
This holds by design: $A_{jt}=I(t\geq\tau_{j})$ and $\tau_{j}$ is randomized independently of individual outcomes.
Assumption 3 (Surrogate-Mediated Missing at Random).
The unobserved outcome provides no additional information about the censoring probability once the surrogate is known (Rubin, 1976):
$P(\Delta_{ijt}=1\mid Y_{ijt},S_{ijt},A_{jt},W_{ijt},t)\;=\;P(\Delta_{ijt}=1\mid S_{ijt},A_{jt},W_{ijt},t).$ (9)
This is structurally encoded in the SCM by the absence of the edge $Y_{ijt}\to\Delta_{ijt}$ (Figure 1, Equation (5)).
Remark 2 (Practical plausibility of Assumption 3).
Assumption 3 is most defensible when missingness is administrative (calendar-driven) rather than informative (health-driven). In the SW-CRT setting, the dominant censoring mechanism — the trial’s fixed end date — is independent of $Y_{ijt}$ conditional on $S_{ijt}$. A strong surrogate candidate satisfies three properties: (i) causal proximity: $S$ lies on or near the pathway $A\to S\to Y$; (ii) temporal availability: $S$ is observed for all individuals regardless of crossover time; and (iii) biological stationarity: $\mathbb{E}[Y\mid S,A,W,t]$ does not change systematically over calendar time. Examples include 3-month viral load predicting 12-month suppression in HIV trials, and discharge-process scores predicting 30-day readmission in quality-improvement SW-CRTs.
Assumption 4 (Support Positivity for the Observed Outcome Regression).
There exists a constant $c>0$ such that, for all $(s,a,w,t)$ in the support of $(S,A,W,t)$ relevant to the target parameter,
$P(\Delta_{ijt}=1\mid S_{ijt}=s,A_{jt}=a,W_{ijt}=w,t)\;\geq\;c.$ (10)
Remark 3 (Support positivity versus inverse-weighting regularity).
Assumption 4 is a support condition for identifying $E[Y\mid S,A,W,t]$ from the observed cohort $\{\Delta=1\}$, not an inverse-weighting regularity condition. IPCW-type estimators are sensitive to small values of $g_{\Delta}$ because $g_{\Delta}^{-1}$ appears directly in the influence function, so that even when Assumption 4 holds with a small constant $c$, the IPCW estimating equation becomes highly variable. The surrogate-bridge parameter, by contrast, does not contain $g_{\Delta}^{-1}$ in the target functional; it requires only that the complete-case outcome regression $E[Y\mid S,A,W,t,\Delta=1]$ be well-defined on the support over which the surrogate-bridge formula integrates.
@@
The marginal treatment probability, averaged over the trial design space, is non-degenerate:
$0\;<\;\mathbb{E}_{t}\bigl[g_{A}(1\mid W,t)\bigr]\;<\;1.$ (11)
Remark 4 (Why Treatment Positivity Holds Marginally Despite Pointwise Violations).
At any fixed time $t$ and cluster $j$, the treatment assignment is deterministic: $A_{jt}=I(t\geq\tau_{j})$, so pointwise positivity $0<P(A_{jt}=1\mid W,t)<1$ fails everywhere. A naive reading would therefore conclude that treatment positivity is structurally violated by the SW-CRT design.
This is resolved by observing that the identification formula (1) marginalizes over the trial’s design space, averaging jointly over both $W$ and $t$. The design-based propensity is known by construction from the randomization schedule. Under a uniform randomization of crossover times:
$g_{A}(1\mid W,t)\;=\;P(\tau_{j}\leq t)\;=\;\frac{t-1}{T-1}.$ (12)
Averaging over the interior steps $t\in\{2,\ldots,T-1\}$ yields $\mathbb{E}_{t}[g_{A}(1\mid W,t)]=1/2$, so Assumption 5 holds. In practice, boundary steps $t=1$ and $t=T$ should be excluded, as $g_{A}$ equals 0 and 1 respectively at these endpoints.
A further consequence is that $g_{A}$ need not be estimated from the data. In Section 4, we plug in the design-based propensity directly, eliminating one source of nuisance estimation error from the second-order remainder $R_{2}$ (Supplement, Section S2).
@@
By Assumptions 1 (Consistency) and 2 (Sequential Randomization), the potential outcome mean satisfies:
$\mathbb{E}[Y(a)\mid S(a),W,t]\;=\;\mathbb{E}[Y\mid S,A=a,W,t].$ (14)
By Assumption 3 (Surrogate-Mediated MAR) and Assumption 4 (Support Positivity), we may condition on the observed cohort ($\Delta=1$):
$\mathbb{E}[Y\mid S,A=a,W,t]\;=\;\mathbb{E}[Y\mid S,A=a,W,t,\Delta=1]\;\equiv\;\bar{Q}_{Y}(S,a,W,t).$ (15)
Integrating $\bar{Q}_{Y}$ over the conditional surrogate distribution:
$\mathbb{E}[Y(a)\mid W=w,t]\;=\;\int\bar{Q}_{Y}(s,a,w,t)\,dP(s\mid a,w,t)\;\equiv\;\bar{Q}_{\mathrm{int}}(a,w,t).$ (16)
Marginalizing over the trial design space yields (1).∎∎

### 15 Mar 2026 v2

# Surrogate-Assisted Targeted Learning for Nested Bridge Functionals under Administrative Censoring

@@ Abstract
We study semiparametric estimation of the average treatment effect when the target functional belongs to the nested bridge class: identification requires integrating an outcome regression over a conditional intermediate distribution rather than inverse-weighting incomplete observations. This structure arises whenever a primary outcome is administratively censored for a non-negligible fraction of units while an earlier surrogate remains broadly available, and it generates a distinctive obstruction for standard first-order estimators: a one-step debiased machine-learning construction leaves a second-order cross-product remainder $R_{SY}$ involving the conditional surrogate law $f_{S}$ that cross-fitting does not eliminate and that has no doubly-robust complement in the efficient influence function.
We develop a surrogate-assisted targeted minimum loss estimator (SA-TMLE) that resolves this obstruction via a two-stage nested fluctuation step, and establish three structural results: (i) under surrogate-mediated missing at random, the censoring mechanism contributes no separate tangent-space component to the efficient influence function; (ii) the standard DML one-step remainder contains $R_{SY}$ as an irreducible second-order term (Proposition 1); and (iii) the two-stage targeting step absorbs $R_{SY}$ into the efficient score without estimating $f_{S}$, yielding an asymptotically linear, doubly robust estimator under a single product-rate condition (Theorem 2). For clustered data, valid inference requires summation rather than averaging of individual influence contributions at the cluster level (Lemma 2); a leave-one-cluster-out jackknife applied to the full estimator restores near-nominal finite-sample coverage where the first-order sandwich interval undercovers.
Stepped-wedge cluster-randomized trials, in which calendar-driven rollout creates near-boundary observation probabilities for late-crossing clusters, provide the sharpest instance of this functional class and serve as the running example. Simulations confirm near-zero bias and jackknife coverage of 0.956–0.968 across all cluster sizes evaluated, against a backdrop where inverse-weighted competitors break down entirely at high censoring rates.
@@ 1 Introduction
A class of causal functionals in modern longitudinal studies shares the following semiparametric structure. The target parameter — an average treatment effect — is identified not by inverse-weighting incomplete observations, but by integrating an outcome regression over the conditional distribution of an intermediate variable. We call this a nested bridge functional: its efficient score has a cross-product structure that creates a specific failure mode for standard first-order estimators, and resolving that failure mode requires a targeting step that goes beyond ordinary debiased machine learning.
The observational setting that generates this structure is one where a primary endpoint requires substantial follow-up to realize, so that only units enrolled early enough have a fully matured observation by the time of analysis. A shorter-term surrogate is observed broadly across the study population regardless of enrollment timing. Rather than inverse-weighting the incomplete observations by $g_{\Delta}^{-1}$ — which becomes unstable whenever $g_{\Delta}$ concentrates near zero — one identifies the average treatment effect by integrating the complete-case outcome regression $\bar{Q}_{Y}(S,a,W,t)\equiv\mathbb{E}[Y\mid S,A=a,W,t,\Delta=1]$ over the treatment-specific conditional distribution of the surrogate:
$\Psi(P_{0})\;=\;\mathbb{E}_{W,t}\Bigl(\mathbb{E}_{S\mid A=1,W,t}\bigl[\bar{Q}_{Y}(S,1,W,t)\bigr]-\mathbb{E}_{S\mid A=0,W,t}\bigl[\bar{Q}_{Y}(S,0,W,t)\bigr]\Bigr).$ (1)
This representation requires only that $\bar{Q}_{Y}$ be estimable on the observed sub-cohort $\{\Delta=1\}$, not that $g_{\Delta}$ be bounded away from zero. The price is a new semiparametric complication: the functional (1) is nested in the sense that the outer integration is over the conditional surrogate law $f_{S}$, and the estimator must marginalize over $f_{S}$ without placing it in a denominator.
The core semiparametric obstruction.
The nested structure of (1) creates a specific failure mode for standard debiased machine-learning (DML) estimators. A natural one-step construction estimates $\bar{Q}_{\mathrm{int}}(a,W,t)\equiv\int\bar{Q}_{Y}(s,a,W,t)\,f_{S}(s\mid a,W,t)\,d\mu(s)$ by plugging in $\hat{\bar{Q}}_{Y}$ and $\hat{f}_{S}$ separately, leaving a second-order cross-product remainder
$R_{SY}\;=\;\sum_{a}\iint\bigl(\hat{\bar{Q}}_{Y}-\bar{Q}^{0}_{Y}\bigr)(s,a,w,t)\,\bigl(\hat{f}_{S}-f^{0}_{S}\bigr)(s\mid a,w,t)\;d\mu(s)\,dP_{0}(w,t).$ (2)
Unlike the standard doubly-robust remainder, $R_{SY}$ has no complement in the efficient influence function: no nuisance parameter whose consistent estimation forces $R_{SY}=0$. DML cross-fitting eliminates first-order empirical-process terms but leaves $R_{SY}$ unchanged, since $R_{SY}$ is a $P_{0}$-expectation of a second-order product. Controlling $R_{SY}=o_{P}(J^{-1/2})$ therefore requires either a rate condition on $\hat{f}_{S}$ or an estimator that avoids placing $\hat{f}_{S}$ in the expansion. The SA-TMLE takes the second path: its two-stage nested fluctuation enforces $N^{-1}\sum D_{S,a}(O_{ijt})=0$, constraining $\hat{\bar{Q}}^{*}_{\mathrm{int}}$ to a propensity-weighted empirical mean of $\hat{\bar{Q}}^{*}_{Y}$ over observed surrogate values, absorbing $R_{SY}$ without ever estimating $f_{S}$.
Motivating instance.
Stepped-wedge cluster-randomized trials (SW-CRTs) provide a sharp instance of this structure. Calendar-determined treatment rollout means late-crossing clusters accumulate few primary-outcome observations before the administrative end date, creating a design-induced near-boundary regime $g_{\Delta}\to 0$ for those clusters. A short-term surrogate observed at crossover is available for all clusters, making the surrogate-bridge framework directly applicable. We use the SW-CRT as the running example throughout; the framework is not tied to this design.
@@
IPCW estimators (Robins et al., 1994; Bang and Robins, 2005) place $g_{\Delta}^{-1}$ directly in the estimating equation, producing severe variance inflation as $g_{\Delta}\to 0$ and potentially failing to admit regular asymptotic approximations near the boundary. Parametric mixed models (Hussey and Hughes, 2007; Hemming and Taljaard, 2020; Hughes et al., 2020) restrict attention to the complete-case sub-cohort and require correct specification of the secular time trend and censoring mechanism (Kasza et al., 2019).
@@
This paper makes three contributions to semiparametric estimation theory, instantiated in the SW-CRT setting.
First, we identify the nested bridge functional class under surrogate-mediated administrative censoring and establish a G-computation representation (Theorem 1) that avoids inverse observation weights entirely. The representation replaces the unstable $g_{\Delta}^{-1}$ with a support positivity condition on the complete-case outcome regression, extending the scope of doubly robust estimation to near-boundary censoring regimes.
Second, we characterize the efficient-score geometry of this functional class (Section 3). Three structural results follow: the censoring mechanism contributes no separate tangent-space component under surrogate-mediated MAR (Lemma 1); the nested cross-product remainder $R_{SY}$ is an irreducible second-order obstruction with no doubly-robust complement (Proposition 1); and valid inference under clustered dependence requires summation rather than averaging of individual influence contributions at the cluster level (Lemma 2).
Third, we construct the SA-TMLE and prove that its two-stage nested fluctuation absorbs $R_{SY}$ without estimating $f_{S}$, achieving $\sqrt{J}$-consistent asymptotic linearity under a single product-rate condition (Theorem 2). For finite-sample inference, a leave-one-cluster-out jackknife applied to the full estimator captures the fluctuation-step variance that the first-order sandwich misses, restoring near-nominal coverage across the cluster-count range evaluated.
@@
Classical AIPW theory (Robins et al., 1994; Bickel et al., 1993) characterizes efficient estimation with missing outcomes but does not address the nested bridge structure or the $R_{SY}$ obstruction. The surrogate outcome literature (Gilbert et al., 2008; Frangakis and Rubin, 2002) addresses the complementary problem of rare (not delayed) primary outcomes; their identification strategies do not extend to the near-boundary censoring regime. Luedtke et al. (2017) construct a semiparametric surrogate-index estimator for i.i.d. data structurally related to ours, but their asymptotic linearity requires a uniform positivity lower bound on $g_{\Delta}$ that the present approach does not need. The DML framework (Chernozhukov et al., 2018) provides product-rate guarantees via sample splitting, but cross-fitting does not eliminate $R_{SY}$; this paper isolates that remainder (Proposition 1) and resolves it via a second targeting step. [^1] See Kennedy (2022) for a review of debiased ML. Cluster-level TMLE (Balzer et al., 2016, 2021) establishes clusterwise influence aggregation for group-randomized trials without the nested bridge structure studied here.
@@
Section 2 defines the observed data structure and establishes non-parametric identification via the surrogate-bridge representation. Section 3 develops the semiparametric theory, including the EIC decomposition, the vanishing censoring-mechanism component (Lemma 1), the cluster-level summation rule (Lemma 2), and Proposition 1 characterizing $R_{SY}$. Section 4 constructs the SA-TMLE. Section 5 establishes asymptotic linearity, double robustness, and the Berry–Esseen coverage bound. Section 6 reports Monte Carlo experiments. Section 7 provides the design-calibrated EPT illustration.
@@ 2 Data Structure and Identification
We now formalize the observed-data structure and identification strategy for the general nested surrogate-bridge problem described in Section 1. The framework applies whenever a delayed primary outcome is administratively censored for a subset of units at analysis time and an earlier surrogate is broadly observed. We instantiate it throughout in the stepped-wedge cluster-randomized setting, which makes the positivity boundary concrete and provides the simulation and illustration target of later sections; the identification theorem and semiparametric theory of Sections 2.23 are stated for the general structure.
@@
Three structural features deserve emphasis. First, $A_{jt}$ is deterministic given $\tau_{j}$; the treatment mechanism is known by design and $g_{A}(1\mid W,t)=(t-1)/(T-1)$ is plugged in directly rather than estimated. Second, the cluster random effect $b_{j}$ enters both $f_{S}$ and $f_{Y}$, inducing the ICC among all observations within cluster $j$ (Remark 1). Third — and most critically — $Y$ does not appear in $f_{\Delta}$ and $U_{\Delta}$ does not appear in $f_{Y}$. This is the structural encoding of Assumption 3: conditional on $S$, the censoring indicator $\Delta$ is independent of the outcome $Y$. In DAG terms, there is no directed edge $Y_{ijt}\to\Delta_{ijt}$ (Figure 1).
@@
The shared $b_{j}$ induces ICC $\sigma^{2}_{b}/(\sigma^{2}_{b}+\sigma^{2}_{\varepsilon})$; inference must use the cluster-robust sandwich of Section 3 and Super Learner cross-validation at the cluster level (Supplement, Section S3).
@@ 2.1 Potential Outcomes and the Causal Target
For $a\in\{0,1\}$, the causal ATE is $\Psi(P_{0})=\mathbb{E}[Y_{ijt}(1)-Y_{ijt}(0)]$, where the expectation is over the marginal distribution of $(W_{ijt},b_{j},t)$.
@@ 2.2 Non-Parametric Identification
$(Y(0),Y(1),S(0),S(1))\perp A_{jt}\mid W_{ijt},t$. This holds by design since $\tau_{j}$ is randomized independently of outcomes.
Assumption 3 (Surrogate-Mediated MAR).
$P(\Delta_{ijt}=1\mid Y_{ijt},S_{ijt},A_{jt},W_{ijt},t)=P(\Delta_{ijt}=1\mid S_{ijt},A_{jt},W_{ijt},t)$ (Rubin, 1976). Assumption 3 is most defensible when censoring is calendar-driven rather than health-driven; strong surrogates satisfy causal proximity ($S$ on the path $A\to S\to Y$), temporal availability, and biological stationarity of $\mathbb{E}[Y\mid S,A,W,t]$.
Assumption 4 (Support Positivity).
$P(\Delta_{ijt}=1\mid S_{ijt}=s,A_{jt}=a,W_{ijt}=w,t)\geq c>0$ on the relevant support. Unlike IPCW, the surrogate-bridge target functional does not contain $g_{\Delta}^{-1}$; this is a support condition for identifying $\mathbb{E}[Y\mid S,A,W,t]$ from $\{\Delta=1\}$, not an inverse-weighting regularity condition.
@@
$0<\mathbb{E}_{t}[g_{A}(1\mid W,t)]<1$. Under uniform crossover randomization $g_{A}(1\mid W,t)=(t-1)/(T-1)$, known by construction, so $g_{A}$ is plugged in directly rather than estimated.
@@
By Assumptions 12, $\mathbb{E}[Y(a)\mid S(a),W,t]=\mathbb{E}[Y\mid S,A=a,W,t]$. By Assumption 3 and 4, conditioning on $\{\Delta=1\}$ is valid: $\mathbb{E}[Y\mid S,A=a,W,t]=\mathbb{E}[Y\mid S,A=a,W,t,\Delta=1]\equiv\bar{Q}_{Y}(S,a,W,t)$. Integrating over $P(S\mid a,W,t)$ and marginalizing over $(W,t)$ yields (1).∎∎